a. An 8%, five-year bond yields 6%. If this yield to maturity remains unchanged, what will be its price one year hence? Assume annual coupon payments and a face value of $100. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Price            $

b. What is the total return to an investor who held the bond over this year? (Do not round intermediate calculations. Enter your answer as a percent rounded to the nearest whole number.)

Total return             %
Answer #1

a)Coupon = 8%*100 = 8
YTM = 6%
Price today = PV of coupons + PV of Par Value = 8*(1-(1+6%)-5)/6% + 100/(1+6%)5 = 108.42

Price after one year = PV of coupons + PV of Par Value = 8*(1-(1+6%)-4)/6% + 100/(1+6%)4 = 106.93

b) Holding period return = (106.93 - 108.42 + 8)/108.42 = 6%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well